OpenStudy (anonymous):

The length of a rectangle is 48in and is dilated to 12 in. What is the scale factor? A.1/4 B.-4 C.4 D.-1/4

jimthompson5910 (jim_thompson5910):

hint: solve 12 = 48k for k

OpenStudy (anonymous):

4 :)

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (anonymous):

oh wait

jimthompson5910 (jim_thompson5910):

12 = 48k 12/48 = k 1/4 = k k = 1/4

jimthompson5910 (jim_thompson5910):

I'm sure you just realized that too

OpenStudy (anonymous):

yeah haha ...

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

Thankyou! :)

jimthompson5910 (jim_thompson5910):

you're welcome